OpenStudy (anonymous):

Medal and Fan. Solve x^2-4x-21=0 by factoring and the zero product property. (ab=0 if a=0 OR b=0)

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to -21 and add to -4 at the same time

OpenStudy (anonymous):

-7 and 3.

jimthompson5910 (jim_thompson5910):

so x^2 - 4x - 21 = 0 turns into (x-7)(x+3) = 0

OpenStudy (anonymous):

Okay. But.. what about the -4x..? I get confused when it comes to factoring.

jimthompson5910 (jim_thompson5910):

I got the -4 from the -4x

jimthompson5910 (jim_thompson5910):

so that's why x^2 - 4x - 21 factors to (x-7)(x+3) -7+3 = -4 -7*3 = -21

OpenStudy (anonymous):

Ohh!! duh. Okay. Continue please (:

jimthompson5910 (jim_thompson5910):

so you have (x-7)(x+3) = 0

jimthompson5910 (jim_thompson5910):

use the zero product property to break up (x-7)(x+3) = 0 into x-7 = 0 or x+3 = 0 x = ??? or x = ???

OpenStudy (anonymous):

So then x=7 or x=-3.. yeah?

jimthompson5910 (jim_thompson5910):

both are correct

OpenStudy (anonymous):

Ahh! thank you, so much easier than I was making it.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Do you have time for one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Okay. I need to know how to find the relative maximum and minimum of a function. I'm given f(x)=2x^3+x^2-11x

jimthompson5910 (jim_thompson5910):

Since you're not in calculus, you'll need to use a graphing calculator to get the answer

jimthompson5910 (jim_thompson5910):

https://www.desmos.com/calculator is a good graphing calculator

OpenStudy (anonymous):

Luckily I'm familiar with that site (: Okay.. so now what?

jimthompson5910 (jim_thompson5910):

relative maximum = highest point in a certain region relative minimum = lowest point in a certain region

jimthompson5910 (jim_thompson5910):

type in 2x^3+x^2-11x and look for the turning points

OpenStudy (anonymous):

Okay.

jimthompson5910 (jim_thompson5910):